Sandy needs 14 lb of metal with 50 % silver. If Sandy combines one metal with 16 % silver, and another with 78 % silver, how much of each metal does Sandy need?
Answer by fractalier(6550)

Call the number of pounds of 16% used, x.
Then the number of pounds of 78% used must be 14-x.
The setup looks like this:
.16x + .78(14-x) = .50(14)
Now solve for x...
.16x + 10.92 - .78x = 7
-.62x = -3.92
x = 6.32 lb of 16%
which means Sandy needs
14-x = 7.68 lb of 78% silver